Find dp/dq If P= q²+3/(q-1)³+(q+1)³


1
Expert's answer
2022-01-12T18:25:23-0500

"p=q²+\frac{3}{(q-1)³}+(q+1)³"

Differentiate both sides of the equation

"\frac{dp}{dq}=\frac{d}{dq}(q²+\frac{3}{(q-1)³}+(q+3)³"

The derivative of p with respect to q is

"p'" .

By the sum rule,

"\implies p'=\frac{d}{dq}(q²)+\frac{d}{dq}\frac{3}{(q-1)³}+\frac{d}{dq}(q+1)³"

Differentiating the right side of the equation

"p'=2q-\frac{9}{(q-1)⁴}+3(q+1)²"


"\implies \frac{dp}{dq}=2q-\frac{9}{(q-1)⁴}+3(q+1)²"